Ali is 8 years older than Fatima. Six years ago, Ali was 5 times as old as Fatima. How old is Fatima? How old will be Ali in 8 years time?
Answer by jorel1380(3719)

a-8=f
a=f+8
a-6=5(f-6)
f+8-6=5f-30
4f=28
f=7
Fatima is 7, Ali is 15
In 8 years, Ali will be 23..